# A Review On Safe Reinforcement Learning Using Lyapunov and Barrier Functions

### v2

@@ 3.4 Barrier Functions
Barrier functions where first introduced in optimization literature, which are added to cost functions to avoid undesirable regions. CBFs have proven to be effective in defining safe sets in the control community [38]. Approximating system models helps alleviate the need for complete prior system knowledge which is required for formulation of CBFs. We briefly cover theoretical aspects of control barrier functions in this review, a comprehensive review of theory and applications of control barrier functions can be found in [38] and barrier functions in [49]. Consider a compact set $\mathcal{C}$, defined as the zero sub-level set of a barrier function $h:\mathcal{X}\mapsto\mathbb{R}$ ($\mathcal{C}={x:h(x)\leq 0}$). From Proposition 1 in [50], if there exists a strictly increasing scalar function $\alpha:\mathbb{R}\mapsto\mathbb{R}$ such that $\alpha(0)=0$ (an extended class-$\mathcal{K}$ function) and
@@
4 Literature Review
4.1 Lyapunov Functions for Reinforcement Learning
@@
The closed-loop system’s stability assurances can be used to ensure safety guarantees. These methods frequently rely on control Lyapunov functions (CLFs) for Lyapunov stability verification presented in Section 3.3. CLFs were formalized for a general nonlinear dynamical system by Sontag [48]. An interesting property that a lot of work in safe RL leverages is the fact that level sets of CLFs are both attractive and invariant. Fig. 4 shows a brief classification of using Lyapunov functions approaches in RL. The earliest work using Lyapunov functions with RL was proposed by Perkins et al. [51], where the authors use Lyapunov functions to design multiple controllers which the control policy can switch between, which is designed using RL (agent decides when to switch between Lyapunov-based control laws). The system is formulated as a MDP and the authors implement state-action-reward-state-action (SARSA) ($\lambda$) RL algorithm with cerebellar model articulation controller (CMAC) function approximator [52] for action-value functions. The use of Lyapunov function ensures safety guarantees for any switching policy and furthermore, empirically showed accelerated learning.
@@
In [53], the authors propose an actor-critic RL algorithm (with NNs as function approximators) and a Lyapunov function as the optimal-value function for a deterministic nonlinear dynamical system. The authors prove asymptotic stability of the proposed algorithm (Section IV in the work). Similarly, in [54], the authors consider a Lyapunov candidate function as the value function and use NNs to approximate the solution for Hamilton-Jacobi-Bellman (HJB) equation. The authors develop a concurrent learning-based implementation of RL and use Lyapunov analysis to guarantee stability for an approximate dynamic programming (ADP) problem. Similar to the previous work, authors in  [55] compute a control Lyapunov-value function (CLVF) using Hamilton-Jacobi reachability (HJR) analysis. HJR analysis computes a reach-avoid set which can be broadly described as set of states from which the system can driven using a control input to a target set while satisfying operational constraints. In [56], the authors use ADP for a nonlinear dynamical system with weight updates using the Lyapunov stability criterion. NNs are used as function approximator for the actor, critic networks and model approximation. The Lyapunov stability criterion based weight updates perform better than gradient descent for parameter variation and disturbance signal effects. Thus, in [53, 54, 56] a similar approach of considering the value function as the candidate Lyapunov function is considered for an ADP formulation, with [56] only differing in the parameter update criterion. Authors in [66] assume the Q-function as a candidate Lyapunov function and present a distributed multi-agent architecture.
@@
Berkenkamp et al. [57] propose use of Gaussian Process (GP) to approximate the system model online and uses region of attraction (states inside region of attraction are safe) to formulate the safety constraints. The region of attraction is selected by taking the largest level set of Lyapunov function. Thus, the policy optimization problem has a constraint in terms of the Lyapunov decrease condition which is given as follows:
@@
where, $p_{n}{x,u}$ is the upper bound on the $v(f(x,u))$, $v$ is the Lyapunov function and dynamics of system is given by $f(x,u)$. $\tau$ is the discretization constant for the discrete dynamics and $\mathcal{X}_{\tau}$ is the discretization of $\mathcal{X}$. $L_{v}$, $L_{f}$ and $L_{\pi}$ are the Lipschitz constants for the Lyapunov function, dynamics and the control policy set $\Pi_{L}$ (Assumption $1$ in the cited work), respectively. Thus, largest level set of $v$ for which all state-action pairs that correspond to discrete states within $\mathcal{X}_{\tau}$ are contained in the set $L^{dec}_{n}$. [79] follows a similar approach as well with the difference being incorporation of the empirical Lyapunov risk term in the reward function. A drawback of these approaches is that the initial policy is assumed to be known and safe for the system. Expected violation of Lyapunov conditions is used as the loss function for critic in [58]. The authors propose a novel physics informed actor-critic architecture where a Zubov function [59] (maximal Lyapunov function) is employed to approximate the critic function which denotes the physics informed part in the algorithm. Overall, the Zubov function (critic) and controller (actor) are iteratively updated ensuring the proposed method satisfies the Lyapunov conditions.
Lyapunov-based Methods        [60, 61, 62]Lyapunov Loss [63, 64, 65, 58]Lyapunov Value Function [53, 54, 66]       [57, 67, 68, 55]Lyapunov-RL-QP [56, 69, 70, 71, 72]       [73, 74]Lyapunov-based reward [75, 76, 77]       [78, 79]
@@
Compared to most prior work based on control affine nonlinear dynamical systems, in [75], authors formulate the safe RL problem as a CMDP. The safety constraints in the CMDP are modeled as constraint cost function (upper bound on cost) using Lyapunov functions, thereby reformulating the value function as a cost-value function. The following problem is solved in this case: Given an initial condition $x_{0}$ and a upper-bound on the expected cumulative constraint cost $d_{0}$, solve
@@
where, $T_{\pi^{*},d}$ is the Bellman operator with respect to optimal policy $\pi^{*}$ and cost constraint function $d$. This framework is extended to DP and RL algorithms (safe policy iteration, safe value iteration, safe Q-learning etc.) in this work. Chow et al. [76] also propose safe policy optimization for a continuous action CMDP formulations and test it on MuJoCo and real-world indoor navigation tasks. Two approaches: $\theta$-projection (constrained optimization which reformulates policy gradient (PG) with a projection of the policy parameters onto a feasible set using Lyapunov functions) and $a$-projection (projection of unconstrained action on the Lyapunov hyperplane) are proposed for solving on and off policy gradient algorithms. Another instance of using CMDP formulation is proposed by authors in [73], where the Lyapunov drift term is used as a constraint for QP problem that projects the RL action to a safe set. This approach is similar to the CLF-QP formulation with CLF constraints are replaced by Lyapunov drift terms.
@@
Perkins and Barto (2002) [51] Switching between Lyapunov control laws SARSA ($\lambda$) Pendulum Scenarios
Vamvoudakis et al. (2015) [53] Lyapunov candidate value function ADP Van der Pol, Power Plant
Kamalapurkar et al. (2016) [54] Lyapunov candidate value function ADP LQT problem
@@
Berkenkamp et al. (2017) [57] Region of attraction based safe set (largest level set of Lyapunov function) ADP Inverted pendulum
@@
Kumar et al. (2017) [56] Lyapunov criterion based weight updates ADP Nonlinear dynamical systems
@@
Chow et al. (2018) [75] Constraints as cost functions using Lyapunov functions Multiple DP and RL algorithms 2D Grid World
@@
Chow et al. (2019) [76] $\theta$-projection, $a$-projection Multiple policy gradient algorithms MuJoCo Simulator
@@
Dong et al. (2020) [77] Lyapunov-based reward shaping DQN, PPO OpenAI Gym, MuJoCo
@@
Huh et al. (2020) [69] Lyapunov-based policy optimization Tabular-Q, DDPG Double integrator, Reacher robot system
@@
Jeddi et al. (2021) [70] Lyapunov-based constraints on policy DQN, PPO, TRPO Grid-world
Chang et al. (2021) [67] Lyapunov-based critics PPO OpenAI Gym, MuJoCo
@@
Westenbroek et al. (2020) [71] Rate of dissipation CLF constraint on policy SAC Double pendulum, Bipedal walking
Westenbroek et al. (2022) [68] Lyapunov candidate value function SAC Cartpole, A1 quadruped
@@
Hejase et al. (2023) [63] Jointly learning CLF and dynamics using DNN. Lyapunov violation penalty on policy loss. DDPG Gymnasium - highway-env
@@
Krishna et al. (2023) [64] Controlled finite-time Lyapunov exponent analysis MPC and DDPG Custom - unsteady fluid flow field
@@
Kumar et al. (2023) [65] Lyapunov drift and penalty considered in loss function for policy update. Multi-agent DDPG PTV Vissim - Traffic simulation
@@
Cao et al. (2023) [73] Safe set evaluated using Lyapunov stability theory. Unsafe actions by RL agent projected on safe set. PPO Custom
@@
Lopez et al. (2024) [55] Decomposition into subsystems and computing control Lyapunov value function (HJ reachability analysis) for each. SAC and PPO Dubins car, lunar lander, drone
@@
Wang et al. (2024) [58] Minimize expected violation of Lyapunov conditions and maximize region of attraction using Zubov’s method. Actor-critic Double Integrator, Van der Pol, inverted pendulum, bicycle tracking
@@
Tesfazgi et al. (2024) [74] Learning CLFs through demonstrations. Inverse RL LASA dataset
@@
Kushwaha et al. (2024) [60] Learning CLFs using ICNNs with Lyapunov penalty term. Soft Actor-Critic Gymnasium
@@
Yao et al. (2025) [66] State value function as the candidate Lyapunov function. Distributed training of Actor-critic agents. Actor-Critic MATLAB
@@
McCutcheon et al. (2025) [61] Lyapunov-based loss term in policy updates. Lyapunov risk term to update neural Lyapunov function. Soft Actor-Critic Gymnasium
@@
Ugurlu et al. (2025) [78] Lyapunov-based penalty term in reward function. PPO OpenAI gym and PX4 Quadrotor
@@
Hao et al. (2025) [62] Lyapunov-based constraints on policy loss function Multiple RL and SRL algorithms Python - Custom Microgrid
@@
Xia et al. (2025) [79] Lyapunov region of attraction (ROA) term in the reward function. To maximize ROA. Model-based RL Python - Custom USV
@@
Long et al. (2025) [72] Learning policy with Lyapunov-based certificates as constraints and penalty terms. PPO, SAC, TD-MPC Gymnasium, DeepMind Control
@@
Reward shaping is another method for accelerating the learning process in RL while preserving optimality of the policy [80]. Authors in [77, 78] explore reward shaping using Lyapunov stability theory. The reward function for RL is modified by adding the additional term:
@@
where, $\lambda$ is a parameter to adjust the influence of reward shaping and $\gamma$ is assumed to be $1$ (discount factor). An asymptotically unbiased optimal greedy policy is also introduced by maximizing Q-function combined with an additional term similar to (16). Deep Q-learning networks (DQN) and proximal policy optimization (PPO) are used to verify the results in MuJoCo and OpenAI gym. Huh et al. [69] propose use of Lyapunov-based shielding to compute a safe policy and also an efficient safe exploration using Lyapunov constraints for a MDP formulation. Safe policies are sampled from the Lyapunov induced safe set which satisfies the probabilistic safety condition up to a threshold similar to the approach using (15). Efficiency in exploration is ensured using experience replay and for safe exploration the authors propose an exploratory policy (different from the safe policy) which drives an agent around the boundary of the safe set. The methodology is tested for tabular Q-learning and deep deterministic policy gradient (DDPG) algorithms for a double integrator and robotic simulations. In [70], the authors propose a similar Lyapunov-based safe policy formulation and a Transformer Neural Network-based approach to account for uncertainty in a CMDP formulation. The Transformer model is used as a memory buffer to capture long-term dependencies for uncertainty, which is used to predict the mean and variance of violating a safety constraint over a horizon. The mean and variance is fed as input to the risk averse action selection method which minimize the joint cost of selecting an action to minimize the mean and variance of violating the safety constraint over a horizon. The proposed algorithm is tested in 2D grid world with static and dynamic obstacles using DQN, PPO and trust region policy optimization (TRPO) algorithms.
A self-learned Lyapunov critic is a function that estimates the region of attraction for a closed loop controlled system. This is achieved by estimating the empirical Lyapunov risk, which provides a measure of how likely the system is to diverge from a given state. Almost Lyapunov conditions [81] solve the problem of incomplete model information by relaxing the Lyapunov conditions, i.e., a set $\Omega$ is introduced which contains the states that violate the Lyapunov conditions, however, as long as each component of $\Omega$ is small enough (bounded) the violations do not effect stability. Chang et al. [67] leverage these properties to propose self-learned almost Lyapunov critics for policy optimization. Authors introduce use of empirical Lyapunov risk ($R_{f,N,\rho}(\theta)$), which acts as the loss function for Lyapunov critic NN and is given as follows:
@@
Stochastic gradient descent is used to minimize the loss function. The proposed algorithm, “Policy optimization with self-learned almost Lyapunov critics algorithm”, follows a structure similar to PPO with the addition of Lyapunov risk minimization for the parameter $\theta$ updates. The algorithm’s performance is compared with Lyapunov actor-critic, soft actor critic and PPO for inverted pendulum, quadrotor control, automobile path-tracking and MuJoCo Hopper. A similar approach is proposed in [61, 62]. Westenbroek et al. [71] propose use of a CLF for an inaccurate model of the system and use the rate of dissipation of the CLF as a constraint (threshold for minimum acceptable rate) on closed-loop controller. The controller parameters are optimized using model-free policy optimization using data sampled from the system. The framework is tested on a double pendulum and under-actuated bipedal robots using soft actor-critic (SAC) algorithm. Westenbroek et al. [68] also propose use of CLF as a value function and show that smaller values for discount factor lead to stabilizing controllers with the choice of CLF as the value function. The proposed algorithm is tested on cartpole and an A1 quadruped using the SAC algorithm with augment value function approach. Authors in [63], train a DDPG agent based on the violations of Lyapunov conditions for nonlinear vehicle dynamics model in a lane-following environment. The two-step approach first learns system dynamics and a CLF using NNs for a current policy and secondly, trains a DDPG agent to maximize reward and minimize violation of Lyapunov conditions using the learned CLF.
Multi-agent RL (MARL) has seen significant theoretical contribution over the last decade, given its wide range of real-world applications. Authors in [64], propose a Lyapunov-based multi-agent DDPG (L-MADDPG) for task scheduling and resource allocation of vehicle edge computing (VEC) assisted vehicle networks. The loss function in MADDPG is augmented with the Lyapunov drift and penalty terms [82].
Barrier-based Methods Barrier-based Loss [83, 84, 85, 86, 87]    [88, 89, 90]Barrier-based Reward [91, 92, 93, 94, 95]Barrier-RL-QP [26, 96, 97, 98, 99, 100]    [101, 102, 103, 104, 105, 106]Bilevel Optimization [107, 108]

### v3

@@ 3.4 Barrier Functions
Barrier functions where first introduced in optimization literature, which are added to cost functions to avoid undesirable regions. CBFs have proven to be effective in defining safe sets in the control community [ames2019control]. Approximating system models helps alleviate the need for complete prior system knowledge which is required for formulation of CBFs. We briefly cover theoretical aspects of control barrier functions in this review, a comprehensive review of theory and applications of control barrier functions can be found in [ames2019control] and barrier functions in [polyak1992modified]. Consider a compact set $\mathcal{C}$, defined as the zero sub-level set of a barrier function $h:\mathcal{X}\mapsto\mathbb{R}$ ($\mathcal{C}={x:h(x)\leq 0}$). From Proposition 1 in [ames2016control], if there exists a strictly increasing scalar function $\alpha:\mathbb{R}\mapsto\mathbb{R}$ such that $\alpha(0)=0$ (an extended class-$\mathcal{K}$ function) and
@@
4 Lyapunov Functions for Reinforcement Learning
@@
The closed-loop system’s stability assurances can be used to ensure safety guarantees. These methods frequently rely on control Lyapunov functions (CLFs) for Lyapunov stability verification presented in Section 3.3. CLFs were formalized for a general nonlinear dynamical system by Sontag [sontag1989universal]. An interesting property that a lot of work in safe RL leverages is the fact that level sets of CLFs are both attractive and invariant. Fig. 4 shows a brief classification of using Lyapunov functions approaches in RL. The earliest work using Lyapunov functions with RL was proposed by Perkins et al. [perkins2002lyapunov], where the authors use Lyapunov functions to design multiple controllers which the control policy can switch between, which is designed using RL (agent decides when to switch between Lyapunov-based control laws). The system is formulated as a MDP and the authors implement state-action-reward-state-action (SARSA) ($\lambda$) RL algorithm with cerebellar model articulation controller (CMAC) function approximators [lane1992theory] for action-value functions. The use of Lyapunov function ensures safety guarantees for any switching policy and furthermore, empirically showed accelerated learning.
@@
In [vamvoudakis2015asymptotically], the authors propose an actor-critic RL algorithm (with NNs as function approximators) and a Lyapunov function as the optimal-value function for a deterministic nonlinear dynamical system. The authors prove asymptotic stability of the proposed algorithm (Section IV in the work). Similarly, in [kamalapurkar2016model], the authors consider a Lyapunov candidate function as the value function and use NNs to approximate the solution for Hamilton-Jacobi-Bellman (HJB) equation. The authors develop a concurrent learning-based implementation of RL and use Lyapunov analysis to guarantee stability for an approximate dynamic programming (ADP) problem. Similar to the previous work, authors in  [lopez2024decomposing] compute a control Lyapunov-value function (CLVF) using Hamilton-Jacobi reachability (HJR) analysis. HJR analysis computes a reach-avoid set which can be broadly described as set of states from which the system can driven using a control input to a target set while satisfying operational constraints. In [kumar2017diagonal], the authors use ADP for a nonlinear dynamical system with weight updates using the Lyapunov stability criterion. NNs are used as function approximator for the actor, critic networks and model approximation. The Lyapunov stability criterion based weight updates perform better than gradient descent for parameter variation and disturbance signal effects. Thus, in [vamvoudakis2015asymptotically, kamalapurkar2016model, kumar2017diagonal] a similar approach of considering the value function as the candidate Lyapunov function is considered for an ADP formulation, with [kumar2017diagonal] only differing in the parameter update criterion. Authors in [yao2025lyapunov] assume the Q-function as a candidate Lyapunov function and present a distributed multi-agent architecture.
@@
Berkenkamp et al. [berkenkamp2017safe] propose use of Gaussian Process (GP) to approximate the system model online and uses region of attraction (states inside region of attraction are safe) to formulate the safety constraints. The region of attraction is selected by taking the largest level set of Lyapunov function. Thus, the policy optimization problem has a constraint in terms of the Lyapunov decrease condition which is given as follows:
@@
where, $p_{n}{x,u}$ is the upper bound on the $v(f(x,u))$, $v$ is the Lyapunov function and dynamics of system is given by $f(x,u)$. $\tau$ is the discretization constant for the discrete dynamics and $\mathcal{X}_{\tau}$ is the discretization of $\mathcal{X}$. $L_{v}$, $L_{f}$ and $L_{\pi}$ are the Lipschitz constants for the Lyapunov function, dynamics and the control policy set $\Pi_{L}$ (Assumption $1$ in the cited work), respectively. Thus, largest level set of $v$ for which all state-action pairs that correspond to discrete states within $\mathcal{X}_{\tau}$ are contained in the set $L^{dec}_{n}$. [xia2024estimating] follows a similar approach as well with the difference being incorporation of the empirical Lyapunov risk term in the reward function. A drawback of these approaches is that the initial policy is assumed to be known and safe for the system. Expected violation of Lyapunov conditions is used as the loss function for critic in [wang2024actor]. The authors propose a novel physics informed actor-critic architecture where a Zubov function [zubov1964methods] (maximal Lyapunov function) is employed to approximate the critic function which denotes the physics informed part in the algorithm. Overall, the Zubov function (critic) and controller (actor) are iteratively updated ensuring the proposed method satisfies the Lyapunov conditions.
Lyapunov-based Methods         [KUSHWAHA202484, mccutcheon2025neural, hao2024lyapunov]Lyapunov Loss [10155918, krishna2023finite, kumar2023task, wang2024actor]Lyapunov Value Function [vamvoudakis2015asymptotically, kamalapurkar2016model, yao2025lyapunov]        [berkenkamp2017safe, chang2021stabilizing, westenbroek2022lyapunov, lopez2024decomposing]Lyapunov-RL-QP [kumar2017diagonal, huh2020safe, jeddi2021lyapunov, westenbroek2020learning, long2025certifying]        [cao2023toward, tesfazgi2024stable]Lyapunov-based reward [chow2018lyapunov, chow2019lyapunov, dong2020principled]        [ugurlu2025lyapunov, xia2024estimating]
@@
Compared to most prior work based on control affine nonlinear dynamical systems, in [chow2018lyapunov], authors formulate the safe RL problem as a CMDP. The safety constraints in the CMDP are modeled as constraint cost function (upper bound on cost) using Lyapunov functions, thereby reformulating the value function as a cost-value function. The following problem is solved in this case: Given an initial condition $x_{0}$ and a upper-bound on the expected cumulative constraint cost $d_{0}$, solve
@@
where, $T_{\pi^{*},d}$ is the Bellman operator with respect to optimal policy $\pi^{*}$ and cost constraint function $d$. This framework is extended to DP and RL algorithms (safe policy iteration, safe value iteration, safe Q-learning etc.) in this work. Chow et al. [chow2019lyapunov] also propose safe policy optimization for a continuous action CMDP formulations and test it on MuJoCo and real-world indoor navigation tasks. Two approaches: $\theta$-projection (constrained optimization which reformulates policy gradient (PG) with a projection of the policy parameters onto a feasible set using Lyapunov functions) and $a$-projection (projection of unconstrained action on the Lyapunov hyperplane) are proposed for solving on and off policy gradient algorithms. Another instance of using CMDP formulation is proposed by authors in [cao2023toward], where the Lyapunov drift term is used as a constraint for QP problem that projects the RL action to a safe set. This approach is similar to the CLF-QP formulation with CLF constraints are replaced by Lyapunov drift terms.
@@
Berkenkamp et al. (2017) [berkenkamp2017safe] Region-of-attraction safe set (largest level set of Lyapunov function) ADP Inverted pendulum
@@
Kumar et al. (2017) [kumar2017diagonal] Lyapunov criterion-based weight updates ADP Nonlinear dynamical systems
@@
Chow et al. (2018) [chow2018lyapunov] Constraints as cost functions using Lyapunov functions Multiple DP/RL algorithms 2D grid world
@@
Chow et al. (2019) [chow2019lyapunov] $\theta$-projection, $a$-projection Multiple policy-gradient algorithms MuJoCo
@@
Dong et al. (2020) [dong2020principled] Lyapunov-based reward shaping DQN, PPO OpenAI Gym, MuJoCo
@@
Huh et al. (2020) [huh2020safe] Lyapunov-based policy optimization Tabular-Q, DDPG Double integrator, Reacher
@@
Jeddi et al. (2021) [jeddi2021lyapunov] Lyapunov-based constraints on policy DQN, PPO, TRPO Grid world
Chang et al. (2021) [chang2021stabilizing] Lyapunov-based critics PPO OpenAI Gym, MuJoCo
@@
Westenbroek et al. (2020) [westenbroek2020learning] Rate-of-dissipation CLF constraint on policy SAC Double pendulum, bipedal walking
Westenbroek et al. (2022) [westenbroek2022lyapunov] Lyapunov candidate value function SAC Cartpole, A1 quadruped
@@
Hejase et al. (2023) [10155918] Jointly learning CLF and dynamics using DNN; Lyapunov violation penalty on policy loss DDPG Gymnasium (highway-env)
@@
Krishna et al. (2023) [krishna2023finite] Controlled finite-time Lyapunov exponent analysis MPC, DDPG Unsteady fluid flow (custom)
@@
Kumar et al. (2023) [kumar2023task] Lyapunov drift/penalty in loss for policy update Multi-agent DDPG PTV Vissim (traffic)
@@
Cao et al. (2023) [cao2023toward] Safe set via Lyapunov stability; unsafe actions projected onto safe set PPO Custom
@@
Lopez et al. (2024) [lopez2024decomposing] Subsystem decomposition and CLF value functions (HJ reachability) SAC, PPO Dubins car, lunar lander, drone
@@
Wang et al. (2024) [wang2024actor] Minimize expected Lyapunov violations; maximize ROA via Zubov’s method Actor–critic Double integrator, Van der Pol, inverted pendulum, bicycle tracking
@@
Tesfazgi et al. (2024) [tesfazgi2024stable] Learning CLFs through demonstrations Inverse RL LASA dataset
@@
Jack et al. (2024) [xx] Learning CLFs using ICNNs with Lyapunov penalty term Soft Actor-Critic Gymnasium
@@
Yao et al. (2025) [yao2025lyapunov] State value function as candidate Lyapunov function; distributed actor–critic training Actor–critic MATLAB
@@
McCutcheon et al. (2025) [mccutcheon2025neural] Lyapunov-based loss term in policy updates; Lyapunov risk to update neural Lyapunov function Soft Actor-Critic Gymnasium
@@
Ugurlu et al. (2025) [ugurlu2025lyapunov] Lyapunov-based penalty term in reward function PPO OpenAI Gym, PX4 quadrotor
@@
Hao et al. (2025) [hao2024lyapunov] Lyapunov-based constraints on policy loss function Multiple RL/SRL algorithms Python (custom microgrid)
@@
Xia et al. (2025) [xia2024estimating] Lyapunov ROA term in reward to maximize ROA Model-based RL Python (custom USV)
@@
Long et al. (2025) [long2025certifying] Lyapunov certificates as constraints and penalty terms PPO, SAC, TD-MPC Gymnasium, DeepMind Control
@@
Reward shaping is another method for accelerating the learning process in RL while preserving optimality of the policy [ng1999policy]. Authors in [dong2020principled, ugurlu2025lyapunov] explore reward shaping using Lyapunov stability theory. The reward function for RL is modified by adding the additional term:
@@
where, $\lambda$ is a parameter to adjust the influence of reward shaping and $\gamma$ is assumed to be $1$ (discount factor). An asymptotically unbiased optimal greedy policy is also introduced by maximizing Q-function combined with an additional term similar to (16). Deep Q-learning networks (DQN) and proximal policy optimization (PPO) are used to verify the results in MuJoCo and OpenAI gym. Huh et al. [huh2020safe] propose use of Lyapunov-based shielding to compute a safe policy and also an efficient safe exploration using Lyapunov constraints for a MDP formulation. Safe policies are sampled from the Lyapunov induced safe set which satisfies the probabilistic safety condition up to a threshold similar to the approach using (15). Efficiency in exploration is ensured using experience replay and for safe exploration the authors propose an exploratory policy (different from the safe policy) which drives an agent around the boundary of the safe set. The methodology is tested for tabular Q-learning and deep deterministic policy gradient (DDPG) algorithms for a double integrator and robotic simulations. In [jeddi2021lyapunov], the authors propose a similar Lyapunov-based safe policy formulation and a Transformer Neural Network-based approach to account for uncertainty in a CMDP formulation. The Transformer model is used as a memory buffer to capture long-term dependencies for uncertainty, which is used to predict the mean and variance of violating a safety constraint over a horizon. The mean and variance is fed as input to the risk averse action selection method which minimize the joint cost of selecting an action to minimize the mean and variance of violating the safety constraint over a horizon. The proposed algorithm is tested in 2D grid world with static and dynamic obstacles using DQN, PPO and trust region policy optimization (TRPO) algorithms.
A self-learned Lyapunov critic is a function that estimates the region of attraction for a closed loop controlled system. This is achieved by estimating the empirical Lyapunov risk, which provides a measure of how likely the system is to diverge from a given state. Almost Lyapunov conditions [liu2020almost] solve the problem of incomplete model information by relaxing the Lyapunov conditions, i.e., a set $\Omega$ is introduced which contains the states that violate the Lyapunov conditions, however, as long as each component of $\Omega$ is small enough (bounded) the violations do not effect stability. Chang et al. [chang2021stabilizing] leverage these properties to propose self-learned almost Lyapunov critics for policy optimization. Authors introduce use of empirical Lyapunov risk ($R_{f,N,\rho}(\theta)$), which acts as the loss function for Lyapunov critic NN and is given as follows:
@@
Stochastic gradient descent is used to minimize the loss function. The proposed algorithm, “Policy optimization with self-learned almost Lyapunov critics algorithm”, follows a structure similar to PPO with the addition of Lyapunov risk minimization for the parameter $\theta$ updates. The algorithm’s performance is compared with Lyapunov actor-critic, soft actor critic and PPO for inverted pendulum, quadrotor control, automobile path-tracking and MuJoCo Hopper. A similar approach is proposed in [mccutcheon2025neural, hao2024lyapunov]. Westenbroek et al. [westenbroek2020learning] propose use of a CLF for an inaccurate model of the system and use the rate of dissipation of the CLF as a constraint (threshold for minimum acceptable rate) on closed-loop controller. The controller parameters are optimized using model-free policy optimization using data sampled from the system. The framework is tested on a double pendulum and under-actuated bipedal robots using soft actor-critic (SAC) algorithm. Westenbroek et al. [westenbroek2022lyapunov] also propose use of CLF as a value function and show that smaller values for discount factor lead to stabilizing controllers with the choice of CLF as the value function. The proposed algorithm is tested on cartpole and an A1 quadruped using the SAC algorithm with augment value function approach. Authors in [10155918], train a DDPG agent based on the violations of Lyapunov conditions for nonlinear vehicle dynamics model in a lane-following environment. The two-step approach first learns system dynamics and a CLF using NNs for a current policy and secondly, trains a DDPG agent to maximize reward and minimize violation of Lyapunov conditions using the learned CLF. A similar approach is followed in [KUSHWAHA202484] with changes to using Koopman operator theory to identify the system dynamics and avoid training an extra DNN.
Multi-agent RL (MARL) has seen significant theoretical contribution over the last decade, given its wide range of real-world applications. Authors in [krishna2023finite], propose a Lyapunov-based multi-agent DDPG (L-MADDPG) for task scheduling and resource allocation of vehicle edge computing (VEC) assisted vehicle networks. The loss function in MADDPG is augmented with the Lyapunov drift and penalty terms [zhuang2019sdn].
Barrier-based Methods Barrier-based Loss [ohnishi2019barrier, zhang2022barrier, zhao2022safe, zhang2024constrained, dey2024p2bpo]      [liu2023safe1, yang2023model, zhao2023barrier]Barrier-based Reward [marvi2021safe, liu2023safe2, ranjan2024barrier, liu2025safe, xie2025certificated]Barrier-RL-QP [cheng2019end, emam2021safe, ma2021model, cai2021safe, yang2024enhancing, dinh2024towards]     [wang2022ensuring, huang2022barrier, song2022safe, cheng2023safe, hou2024safe, kalaria2024disturbance]Bilevel Optimization [wang2023enforcing, sabouni2024reinforcement]